# Inference-Time Alignment for Diffusion Models via Doob’s Matching

### v1

@@ C.3  Auxilary lemmas for the oracle inequality
(C.15) $=-\underbrace{\int_{{\mathbb{R}}^{d}}\nabla\cdot(\nabla h_{t}^{*}(\mathbf{x})p_{T-t}(\mathbf{x})v_{t}(\mathbf{x}))\,\mathrm{d}\mathbf{x}}_{\text{(i)}}+\underbrace{\int_{{\mathbb{R}}^{d}}\nabla\cdot(\nabla h_{t}^{*}(\mathbf{x})p_{t}(\mathbf{x}))v_{t}(\mathbf{x})\,\mathrm{d}\mathbf{x}}_{\text{(ii)}}.$
Before addressing the term (i) in (C.15), we first restrict this integral on a $d$-dimensional ball $B({\mathbf{0}},R)\coloneq\{\mathbf{x}:\|\mathbf{x}\|_{2}\leq R\}$:
(C.16) $\int_{B({\mathbf{0}},R)}\nabla\cdot(\nabla h_{t}^{*}(\mathbf{x})p_{T-t}(\mathbf{x})v_{t}(\mathbf{x}))\,\mathrm{d}\mathbf{x}=\int_{\partial B({\mathbf{0}},R)}\langle\nabla h_{t}^{*}(\mathbf{x}),\mathbf{n}\rangle p_{T-t}(\mathbf{x})v_{t}(\mathbf{x})\,\mathrm{d}S,$
where we used Gauss’s divergence theorem. Here $\partial B({\mathbf{0}},R)$ is the boundary of $B({\mathbf{0}},R)$, and $\mathbf{n}$ is the outward-pointing unit normal vector on $\partial B({\mathbf{0}},R)$. Furthermore,
$\Big|\int_{\partial B({\mathbf{0}},R)}\langle\nabla h_{t}^{*}(\mathbf{x}),\mathbf{n}\rangle p_{T-t}(\mathbf{x})v_{t}(\mathbf{x})\,\mathrm{d}S\Big|$
$\leq\int_{\partial B({\mathbf{0}},R)}|\langle\nabla h_{t}^{*}(\mathbf{x}),\mathbf{n}\rangle p_{T-t}(\mathbf{x})v_{t}(\mathbf{x})|\,\mathrm{d}S$
$\leq\int_{\partial B({\mathbf{0}},R)}\sup_{\mathbf{x}\in B({\mathbf{0}},R)}\|\nabla h_{t}^{*}(\mathbf{x})\|_{2}\sup_{\mathbf{x}\in B({\mathbf{0}},R)}|v_{t}(\mathbf{x})|p_{T-t}(\mathbf{x})\,\mathrm{d}S$
$\leq\frac{2\sqrt{d}\bar{B}}{(2\pi)^{d/2}\sigma_{t}^{d+2}}\sup_{\mathbf{x}\in B({\mathbf{0}},R)}|v_{t}(\mathbf{x})|\exp\Big(-\frac{(R-1)^{2}}{2\sigma_{T-t}^{2}}\Big)\int_{\partial B({\mathbf{0}},R)}\,\mathrm{d}S$
$=\frac{2\sqrt{d}A_{d-1}\bar{B}}{(2\pi)^{d/2}\sigma_{T-t}^{d+2}}\sup_{\mathbf{x}\in B({\mathbf{0}},R)}|v_{t}(\mathbf{x})|\exp\Big(-\frac{(R-1)^{2}}{2\sigma_{T-t}^{2}}\Big)R^{d-1},$
where $A_{d-1}$ is the surface area of the unit ball in $d$-dimensional space, the first inequality is due to Jensen’s inequality, the second inequality holds from $|\langle\nabla h_{t}^{*}(\mathbf{x}),\mathbf{n}\rangle|\leq\|\nabla h_{t}^{*}(\mathbf{x})\|_{2}\|\mathbf{n}\|_{2}=\|\nabla h_{t}^{*}(\mathbf{x})\|_{2}$, and the third inequality involves Lemma C.4, and the fact that
$p_{T-t}(\mathbf{x})$ $\leq\sup_{\mathbf{x}_{0}\in B({\mathbf{0}},1)}\varphi_{d}(\mathbf{x};\mu_{T-t}\mathbf{x}_{0},\sigma_{T-t}^{2}\mathbf{I}_{d})\int p_{0}(\mathbf{x}_{0})\,\mathrm{d}\mathbf{x}_{0}$
$\leq\frac{1}{(2\pi)^{d/2}\sigma_{T-t}^{d}}\exp\Big(-\frac{(\|\mathbf{x}\|_{2}-1)^{2}}{2\sigma_{T-t}^{2}}\Big).$
Here we used Assumption 1. As a result,
$\lim_{R\rightarrow\infty}\Big|\int_{\partial B({\mathbf{0}},R)}\langle\nabla h_{t}^{*}(\mathbf{x}),\mathbf{n}\rangle p_{T-t}(\mathbf{x})v_{t}(\mathbf{x})\,\mathrm{d}S\Big|=0.$
Combining this equality with (C.16) yields
(C.17) $\int_{{\mathbb{R}}^{d}}\nabla\cdot(\nabla h_{t}^{*}(\mathbf{x})p_{T-t}(\mathbf{x})v_{t}(\mathbf{x}))\,\mathrm{d}\mathbf{x}\,\mathrm{d}\mathbf{x}=\lim_{R\rightarrow\infty}\int_{B({\mathbf{0}},R)}\nabla\cdot(\nabla h_{t}^{*}(\mathbf{x})p_{T-t}(\mathbf{x})v_{t}(\mathbf{x}))\,\mathrm{d}\mathbf{x}=0.$
For the term (ii) in (C.15),
@@
$\int_{{\mathbb{R}}^{d}}\nabla\cdot(\nabla h_{t}^{*}(\mathbf{x})p_{T-t}(\mathbf{x}))v_{t}(\mathbf{x})\,\mathrm{d}\mathbf{x}$
$=\int_{{\mathbb{R}}^{d}}(\Delta h_{t}^{*}(\mathbf{x})p_{T-t}(\mathbf{x})+\nabla h_{t}^{*}(\mathbf{x})\cdot\nabla p_{T-t}(\mathbf{x}))v_{t}(\mathbf{x})\,\mathrm{d}\mathbf{x}$
(C.18) $=(\Delta h_{t}^{*}+\nabla h_{t}^{*}\cdot\nabla\log p_{T-t},v_{t})_{L^{2}(p_{T-t})}.$
Substituting (C.17) and (C.18) into (C.15) yields
(C.19) $(\nabla h_{t}^{*},\nabla v_{t})_{L^{2}(p_{T-t})}=(\Delta h_{t}^{*}+\nabla h_{t}^{*}\cdot\nabla\log p_{T-t},v_{t})_{L^{2}(p_{T-t})}.$
Then using Cauchy-Schwarz inequality and triangular inequality implies
$|(\nabla h_{t}^{*},\nabla v_{t})_{L^{2}(p_{T-t})}|$
$\leq\big\{\|\Delta h_{t}^{*}\|_{L^{2}(p_{T-t})}+\|\nabla h_{t}^{*}\cdot\nabla\log p_{T-t}\|_{L^{2}(p_{T-t})}\big\}\|v_{t}\|_{L^{2}(p_{T-t})}$
$\leq\frac{1}{2\xi}\|\Delta h_{t}^{*}\|_{L^{2}(p_{T-t})}^{2}+\frac{1}{2\xi}\|\nabla h_{t}^{*}\cdot\nabla\log p_{T-t}\|_{L^{2}(p_{T-t})}^{2}+\xi\|v_{t}\|_{L^{2}(p_{T-t})}^{2}$
$\leq\frac{1}{2\xi}\|\Delta h_{t}^{*}\|_{L^{2}(p_{T-t})}^{2}+\frac{1}{2\xi}\|\nabla h_{t}^{*}\|_{L^{2}(p_{T-t})}^{2}\|\nabla\log p_{T-t}\|_{L^{2}(p_{T-t})}^{2}+\xi\|v_{t}\|_{L^{2}(p_{T-t})}^{2}$
(C.20) $\leq\frac{18}{\xi}\frac{d^{2}\bar{B}^{2}}{\sigma_{t}^{8}}+\frac{2}{\xi}\frac{d\bar{B}^{2}}{\sigma_{t}^{4}}\|\nabla\log p_{T-t}\|_{L^{2}(p_{T-t})}^{2}+\xi\|v_{t}\|_{L^{2}(p_{T-t})}^{2},$
where the first inequality holds from (C.19) and Cauchy-Schwarz inequality, the second inequality is due to the inequality $ab\leq\frac{1}{2\xi}a+\frac{\xi}{2}b$, and the last inequality invokes Lemma C.4 and C.5. It remains to estimate the $L^{2}(p_{T-t})$-norm of the score $\nabla\log p_{T-t}$. Indeed,
$\|\nabla\log p_{T-t}\|_{L^{2}(p_{T-t})}^{2}$
(C.21) $=\frac{2}{\sigma_{T-t}^{4}}\Big\{\mu_{T-t}^{2}{\mathbb{E}}\big[\|\mathbf{X}_{0}\|_{2}^{2}\big]+\sigma_{T-t}^{2}{\mathbb{E}}\big[\|{\bm{\varepsilon}}\|_{2}^{2}\big]\Big\}+\frac{2d\mu_{T-t}^{2}}{\sigma_{T-t}^{4}}\leq\frac{6d}{\sigma_{T-t}^{4}},$
where the first equality is owing to Lemma C.2, the first inequality used Assumption 1, and the second and third equalities hold from $\mathbf{X}_{T-t}\stackrel{{\scriptstyle\,\mathrm{d}}}{{=}}\mu_{T-t}\mathbf{X}_{0}+\sigma_{T-t}{\bm{\varepsilon}}$ where $\mathbf{X}_{0}$ is independent ${\bm{\varepsilon}}\sim{\mathcal{N}}({\mathbf{0}},\mathbf{I}_{d})$. The last inequality also use Assumption 1. Substituting (C.21) into (C.20) completes the proof. ∎
@@
According to the standard techniques of symmetrization [Mohri et al., 2018, Theorem 3.3], we have the following generalization bounds.
Lemma C.7.
@@
${\mathbb{E}}\Big[\|\widehat{h}_{t}^{\lambda}-h_{t}^{*}\|_{L^{2}(p_{T-t})}^{2}\Big]\leq{\mathbb{E}}\Big[\frac{1}{n}\sum_{i=1}^{n}(\widehat{h}_{t}^{\lambda}(\mathbf{X}_{T-t}^{i})-h_{t}^{*}(\mathbf{X}_{T-t}^{i}))^{2}\Big]+8\bar{B}{\mathfrak{R}}_{n}({\mathscr{H}}_{t}),$
@@
Proof of Lemma C.7.
Let $\mathbf{X}_{T-t}^{1,\prime},\ldots,\mathbf{X}_{T-t}^{n,\prime}$ be independent copies of $\mathbf{X}_{T-t}^{1},\ldots,\mathbf{X}_{T-t}^{n}$. Let $\varepsilon_{1},\ldots,\varepsilon_{n}$ be a set of i.i.d. Rademacher variables, which are independent of $\mathbf{X}_{T-t}^{1,\prime},\ldots,\mathbf{X}_{T-t}^{n,\prime}$ and $\mathbf{X}_{T-t}^{1},\ldots,\mathbf{X}_{T-t}^{n}$. Since $\widehat{h}_{t}^{\lambda}\in{\mathscr{H}}_{t}$, it follows that
${\mathbb{E}}\Bigg[\|\widehat{h}_{t}^{\lambda}-h_{t}^{*}\|_{L^{2}(p_{T-t})}^{2}-\frac{1}{n}\sum_{i=1}^{n}(\widehat{h}_{t}^{\lambda}(\mathbf{X}_{T-t}^{i})-h_{t}^{*}(\mathbf{X}_{T-t}^{i}))^{2}\Bigg]$
$={\mathbb{E}}\Bigg[{\mathbb{E}}\big[(\widehat{h}_{t}^{\lambda}(\mathbf{X}_{T-t})-h_{t}^{*}(\mathbf{X}_{T-t}))^{2}|\mathbf{X}_{T-t}^{1:n}\big]-\frac{1}{n}\sum_{i=1}^{n}(\widehat{h}_{t}^{\lambda}(\mathbf{X}_{T-t}^{i})-h_{t}^{*}(\mathbf{X}_{T-t}^{i}))^{2}\Bigg]$
@@
(C.22) $=2{\mathbb{E}}\Bigg[\sup_{h_{t}\in{\mathscr{H}}_{t}}\frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}(h_{t}(\mathbf{X}_{T-t}^{i})-h_{t}^{*}(\mathbf{X}_{T-t}^{i}))^{2}\Bigg]\leq 8\bar{B}{\mathfrak{R}}_{n}({\mathscr{H}}_{t}),$
where the second inequality holds from Jensen’s inequality, and last inequality is due to Ledoux-Talagrand contraction inequality [Mohri et al., 2018, Lemma 5.7] and Lemma C.1. This completes the proof. ∎
By a similar argument as Lemma C.7, we have the following generalization bounds.
Lemma C.8.
@@
${\mathbb{E}}\Big[\|\nabla\widehat{h}_{t}^{\lambda}-\nabla h_{t}^{*}\|_{L^{2}(p_{T-t})}^{2}\Big]\leq{\mathbb{E}}\Big[\frac{1}{n}\sum_{i=1}^{n}(\nabla\widehat{h}_{t}^{\lambda}(\mathbf{X}_{T-t}^{i})-\nabla h_{t}^{*}(\mathbf{X}_{T-t}^{i}))^{2}\Big]+\frac{16\bar{B}}{\sigma_{T-t}^{2}}{\mathfrak{R}}_{n}(\nabla{\mathscr{H}}_{t}),$
where the expectation is taken with respect to $\mathbf{X}_{T-t}^{1},\ldots,\mathbf{X}_{T-t}^{i}\sim^{\mathrm{i.i.d.}}p_{T-t}$, and ${\mathfrak{R}}_{n}(\nabla{\mathscr{H}}_{t})\coloneq\sum_{k=1}^{d}{\mathfrak{R}}_{n}(D_{k}{\mathscr{H}}_{t})$.
Lemma C.9.
Suppose Assumptions 1 and 2 hold. Then
${\mathbb{E}}\Big[\frac{1}{n}\sum_{i=1}^{n}\nabla h_{t}^{*}(\mathbf{X}_{T-t}^{i})\cdot\nabla\widehat{h}_{t}^{\lambda}(\mathbf{X}_{T-t}^{i})\Big]\leq{\mathbb{E}}\Big[(\nabla h_{t}^{*},\nabla\widehat{h}_{t}^{\lambda})_{L^{2}(p_{T-t})}\Big]+\frac{8\bar{B}}{\sigma_{T-t}^{2}}{\mathfrak{R}}_{n}(\nabla{\mathscr{H}}_{t}),$
where the expectation is taken with respect to $\mathbf{X}_{T-t}^{1},\ldots,\mathbf{X}_{T-t}^{i}\sim^{\mathrm{i.i.d.}}p_{T-t}$, and ${\mathfrak{R}}_{n}(\nabla{\mathscr{H}}_{t})\coloneq\sum_{k=1}^{d}{\mathfrak{R}}_{n}(D_{k}{\mathscr{H}}_{t})$.
@@
${\mathbb{E}}\Big[\frac{1}{n}\sum_{i=1}^{n}\widehat{h}_{t}^{\lambda}(\mathbf{X}_{T-t}^{i})(w(\mathbf{X}_{0}^{i})-{\mathbb{E}}[w(\mathbf{X}_{0}^{i})|\mathbf{X}_{T-t}^{i}])\Big|\mathbf{X}_{T-t}^{1:n}\Big]\leq 2\bar{B}{\mathfrak{R}}({\mathscr{H}}_{t}|\mathbf{X}_{T-t}^{1:n}),$
@@
Proof of Lemma C.11.
Define a sequence of random variables
$\xi_{i}=w(\mathbf{X}_{0}^{i})-{\mathbb{E}}[w(\mathbf{X}_{0}^{i})|\mathbf{X}_{T-t}^{i}].$
It is apparent that ${\mathbb{E}}[\xi_{i}|\mathbf{X}_{T-t}^{i}]=0$, and $|\xi_{i}|\leq\bar{B}$. Using Lemma C.10 yields
${\mathbb{E}}\Big[\frac{1}{n}\sum_{i=1}^{n}\xi_{i}\widehat{h}_{t}^{\lambda}(\mathbf{X}_{T-t}^{i})\Big|\mathbf{X}_{T-t}^{1:n}\Big]$ $\leq 2\bar{B}{\mathbb{E}}\Big[\frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}\widehat{h}_{t}^{\lambda}(\mathbf{X}_{T-t}^{i})\Big|\mathbf{X}_{T-t}^{1:n}\Big]$
$\leq 2\bar{B}{\mathbb{E}}\Big[\sup_{h_{t}\in{\mathscr{H}}_{t}}\frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}h_{t}(\mathbf{X}_{T-t}^{i})\Big|\mathbf{X}_{T-t}^{1:n}\Big]$
$=2\bar{B}{\mathfrak{R}}({\mathscr{H}}_{t}|\mathbf{X}_{T-t}^{1:n}),$

### v2

@@ D.1  Oracle inequality of variationally stable Doob’s matching
$\quad+\lambda\underbrace{{\mathbb{E}}\Bigg[\sup_{h_{t}\in{\mathscr{H}}_{t}}{\mathbb{E}}\big[\|\nabla h_{t}(\mathbf{X}_{T-t})\|_{2}^{2}\big]-\frac{1}{n}\sum_{i=1}^{n}\|\nabla h_{t}(\mathbf{X}_{T-t}^{i})\|_{2}^{2}\Bigg]}_{\text{(G2)}}.$
We start from the term (G1) in (D.4). First, recall Proposition 4.1:
(D.5) ${\mathbb{E}}\big[(h_{t}(\mathbf{X}_{T-t})-w(\mathbf{X}_{0}))^{2}\big]={\mathbb{E}}\big[(h_{t}(\mathbf{X}_{T-t})-h_{t}^{*}(\mathbf{X}_{T-t}))^{2}\big]+V_{t}.$
For the empirical counterpart, it is straightforward that
$-\frac{1}{n}\sum_{i=1}^{n}(h_{t}(\mathbf{X}_{T-t}^{i})-w(\mathbf{X}_{0}^{i}))^{2}$
$=-\frac{1}{n}\sum_{i=1}^{n}(h_{t}(\mathbf{X}_{T-t}^{i})-h_{t}^{*}(\mathbf{X}_{T-t}^{i})+{\mathbb{E}}[w(\mathbf{X}_{0}^{i})|\mathbf{X}_{T-t}^{i}]-w(\mathbf{X}_{0}^{i}))^{2}$
$=-\frac{1}{n}\sum_{i=1}^{n}(h_{t}(\mathbf{X}_{T-t}^{i})-h_{t}^{*}(\mathbf{X}_{T-t}^{i}))^{2}-\frac{1}{n}\sum_{i=1}^{n}({\mathbb{E}}[w(\mathbf{X}_{0}^{i})|\mathbf{X}_{T-t}^{i}]-w(\mathbf{X}_{0}^{i}))^{2}$
$\quad-\frac{2}{n}\sum_{i=1}^{n}(h_{t}(\mathbf{X}_{T-t}^{i})-h_{t}^{*}(\mathbf{X}_{T-t}^{i}))({\mathbb{E}}[w(\mathbf{X}_{0}^{i})|\mathbf{X}_{T-t}^{i}]-w(\mathbf{X}_{0}^{i})),$
where the first equality invokes the definition of the Doob’s $h$-function $h_{t}^{*}$ in (3.10). Taking expectation with respect to $\{(\mathbf{X}_{0}^{i},\mathbf{X}_{T-t}^{i})\}_{i=1}^{n}$ yields
${\mathbb{E}}\Bigg[\sup_{h_{t}\in{\mathscr{H}}_{t}}{\mathbb{E}}\big[(h_{t}(\mathbf{X}_{T-t})-w(\mathbf{X}_{0}))^{2}\big]-\frac{1}{n}\sum_{i=1}^{n}(h_{t}(\mathbf{X}_{T-t}^{i})-w(\mathbf{X}_{0}^{i}))^{2}\Bigg]$
$={\mathbb{E}}\Bigg[\sup_{h_{t}\in{\mathscr{H}}_{t}}{\mathbb{E}}\big[(h_{t}(\mathbf{X}_{T-t})-h_{t}^{*}(\mathbf{X}_{T-t}))^{2}\big]-\frac{1}{n}\sum_{i=1}^{n}(h_{t}(\mathbf{X}_{T-t}^{i})-h_{t}^{*}(\mathbf{X}_{T-t}^{i}))^{2}\Bigg]$
$\quad+2{\mathbb{E}}\Bigg[\frac{1}{n}\sum_{i=1}^{n}h_{t}(\mathbf{X}_{T-t}^{i})({\mathbb{E}}[w(\mathbf{X}_{0}^{i})|\mathbf{X}_{T-t}^{i}]-w(\mathbf{X}_{0}^{i}))\Bigg]$
(D.6) $\leq 64\bar{B}^{2}\Big(\frac{\mathrm{VCdim}({\mathscr{H}}_{t})}{n\log^{-1}n}\Big)^{\frac{1}{2}}+16\bar{B}^{2}\Big(\frac{\mathrm{VCdim}({\mathscr{H}}_{t})}{n\log^{-1}n}\Big)^{\frac{1}{2}}=80\bar{B}^{2}\Big(\frac{\mathrm{VCdim}({\mathscr{H}}_{t})}{n\log^{-1}n}\Big)^{\frac{1}{2}},$
where the equality invokes (D.5), and the inequality holds from Lemmas D.2 and D.5. For the term (G2) in (D.4), using Lemma D.3 implies
@@
(D.7) ${\mathbb{E}}\Bigg[\sup_{h_{t}\in{\mathscr{H}}_{t}}{\mathbb{E}}\big[\|\nabla h_{t}(\mathbf{X}_{T-t})\|_{2}^{2}\big]-\frac{1}{n}\sum_{i=1}^{n}\|\nabla h_{t}(\mathbf{X}_{T-t}^{i})\|_{2}^{2}\Bigg]\leq\frac{8d\bar{B}^{2}}{\sigma_{T-t}^{4}}\Big(\frac{\mathrm{VCdim}(\nabla{\mathscr{H}}_{t})}{n\log^{-1}n}\Big)^{\frac{1}{2}}.$
Substituting (D.6) and (D.7) into (D.4) yields a generalization error bound:
(D.8) ${\mathbb{E}}\Bigg[\sup_{h_{t}\in{\mathscr{H}}_{t}}{\mathcal{J}}_{t}^{\lambda}(h_{t})-\widehat{{\mathcal{J}}}_{t}^{\lambda}(h_{t})\Bigg]\leq 80\bar{B}^{2}\Big(\frac{\mathrm{VCdim}({\mathscr{H}}_{t})}{n\log^{-1}n}\Big)^{\frac{1}{2}}+\frac{8d\bar{B}^{2}}{\sigma_{T-t}^{4}}\Big(\frac{\mathrm{VCdim}(\nabla{\mathscr{H}}_{t})}{n\log^{-1}n}\Big)^{\frac{1}{2}}.$
Step 2. Approximation error in (D.3). According to the proof of Proposition 4.3, we have
(D.9) ${\mathcal{J}}_{t}^{\lambda}(h_{t})-{\mathcal{J}}_{t}^{\lambda}(h_{t}^{\lambda})=\|h_{t}-h_{t}^{\lambda}\|_{L^{2}(p_{T-t})}^{2}+\lambda\|\nabla h_{t}-\nabla h_{t}^{\lambda}\|_{L^{2}(p_{T-t})}^{2}.$
Step 3. Conclusion. Substituting (D.8) and (D.9) into (D.3) completes the proof. ∎
D.2  Auxilary lemmas for the oracle inequality
@@
According to the standard techniques of symmetrization [Mohri et al., 2018, Theorem 3.3], we have the following generalization bounds. We introduce the concept of Rademacher complexity [Bartlett and Mendelson, 2002, Mohri et al., 2018], which is crucial for analyzing the generalization error.
Definition 3 (Rademacher complexity).
${\mathfrak{R}}({\mathscr{H}}|\mathbf{X}^{1:n})\coloneq{\mathbb{E}}\Big[\sup_{h\in{\mathscr{H}}}\frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}h(\mathbf{X}^{i})\Big|\mathbf{X}^{1:n}\Big],$
Lemma D.2.
@@
${\mathbb{E}}\Bigg[\sup_{h_{t}\in{\mathscr{H}}_{t}}\|h_{t}-h_{t}^{*}\|_{L^{2}(p_{T-t})}^{2}-\frac{1}{n}\sum_{i=1}^{n}(h_{t}(\mathbf{X}_{T-t}^{i})-h_{t}^{*}(\mathbf{X}_{T-t}^{i}))^{2}\Bigg]\leq 64\bar{B}^{2}\Big(\frac{\mathrm{VCdim}({\mathscr{H}}_{t})}{n\log^{-1}n}\Big)^{\frac{1}{2}},$
@@
Proof of Lemma D.2.
Let $\mathbf{X}_{T-t}^{1,\prime},\ldots,\mathbf{X}_{T-t}^{n,\prime}$ be independent copies of $\mathbf{X}_{T-t}^{1},\ldots,\mathbf{X}_{T-t}^{n}$. Let $\varepsilon_{1},\ldots,\varepsilon_{n}$ be a set of i.i.d. Rademacher variables, which are independent of $\mathbf{X}_{T-t}^{1,\prime},\ldots,\mathbf{X}_{T-t}^{n,\prime}$ and $\mathbf{X}_{T-t}^{1},\ldots,\mathbf{X}_{T-t}^{n}$. It follows that
${\mathbb{E}}\Bigg[\sup_{h_{t}\in{\mathscr{H}}_{t}}\|h_{t}-h_{t}^{*}\|_{L^{2}(p_{T-t})}^{2}-\frac{1}{n}\sum_{i=1}^{n}(h_{t}(\mathbf{X}_{T-t}^{i})-h_{t}^{*}(\mathbf{X}_{T-t}^{i}))^{2}\Bigg]$
@@
(D.10) $=2{\mathbb{E}}\Bigg[\sup_{h_{t}\in{\mathscr{H}}_{t}}\frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}(h_{t}(\mathbf{X}_{T-t}^{i})-h_{t}^{*}(\mathbf{X}_{T-t}^{i}))^{2}\Bigg]\leq 8\bar{B}{\mathfrak{R}}_{n}({\mathscr{H}}_{t}),$
where the second inequality holds from Jensen’s inequality, and last inequality is due to Ledoux-Talagrand contraction inequality [Mohri et al., 2018, Lemma 5.7] and Lemma C.1.
It remains to bound the Rademacher complexity ${\mathfrak{R}}_{n}({\mathscr{H}}_{t})$ in (D.10). Let $\delta>0$ and ${\mathscr{H}}_{t}^{\delta}$ be an $L^{\infty}(\mathbf{X}_{T-t}^{1:n})$ $\delta$-cover of ${\mathscr{H}}_{t}$ satisfying $|{\mathscr{H}}_{t}^{\delta}|=N(\delta,{\mathscr{H}}_{t},L^{\infty}(\mathbf{X}_{T-t}^{1:n}))$. Then for any $h_{t}\in{\mathscr{H}}_{t}$, there exists $h_{t}^{\delta}\in{\mathscr{H}}_{t}^{\delta}$ such that
As a consequence,
${\mathfrak{R}}({\mathscr{H}}_{t}\mid\mathbf{X}_{T-t}^{1:n})$ $={\mathbb{E}}\Bigg[\sup_{h_{t}\in{\mathscr{H}}_{t}}\frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}h_{t}(\mathbf{X}_{t}^{i})\mid\mathbf{X}_{T-t}^{1:n}\Bigg]$
$\leq{\mathbb{E}}\Bigg[\sup_{h_{t}^{\delta}\in{\mathscr{H}}_{t}^{\delta}}\frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}h_{t}^{\delta}(\mathbf{X}_{t}^{i})\mid\mathbf{X}_{T-t}^{1:n}\Bigg]+\delta$
$\leq\bar{B}\Big(\frac{2\log|{\mathscr{H}}_{t}^{\delta}|}{n}\Big)^{\frac{1}{2}}+\delta$
$=\bar{B}\Big(\frac{2\log N(\delta,{\mathscr{H}}_{t},L^{\infty}(\mathbf{X}_{T-t}^{1:n}))}{n}\Big)^{\frac{1}{2}}+\delta,$
where $\varepsilon_{1},\ldots,\varepsilon_{n}$ are a sequence of i.i.d. Rademacher variables, the second inequality follows from Massart’s lemma [Mohri et al., 2018, Theorem 3.7], and the equality is due to the definition of ${\mathscr{H}}_{t}^{\delta}$. Then setting $\delta=\bar{B}/\sqrt{n}$ yields
(D.11) ${\mathfrak{R}}({\mathscr{H}}_{t}\mid\mathbf{X}_{T-t}^{1:n})\leq\bar{B}\Big(\frac{2\log N(B/\sqrt{n},{\mathscr{H}}_{t},L^{\infty}(\mathbf{X}_{T-t}^{1:n}))}{n}\Big)^{\frac{1}{2}}\leq 8\bar{B}\Big(\frac{\mathrm{VCdim}({\mathscr{H}}_{t})}{n\log^{-1}n}\Big)^{\frac{1}{2}},$
where the last inequality holds from Anthony et al. [1999, Theorem 12.2]. Substituting (D.11) into (D.10) completes the proof. ∎
By a similar argument as Lemma D.2, we have the following generalization bounds for the gradient term.
Lemma D.3.
@@
${\mathbb{E}}\Bigg[\sup_{h_{t}\in{\mathscr{H}}_{t}}\|\nabla h_{t}\|_{L^{2}(p_{T-t})}^{2}-\frac{1}{n}\sum_{i=1}^{n}\|\nabla h_{t}(\mathbf{X}_{T-t}^{i})\|_{2}^{2}\Bigg]\leq\frac{8d\bar{B}^{2}}{\sigma_{T-t}^{4}}\Big(\frac{\mathrm{VCdim}(\nabla{\mathscr{H}}_{t})}{n\log^{-1}n}\Big)^{\frac{1}{2}},$
Proof of Lemma D.3.
It is straightforward that
${\mathbb{E}}\Bigg[\sup_{h_{t}\in{\mathscr{H}}_{t}}\|\nabla h_{t}\|_{L^{2}(p_{T-t})}^{2}-\frac{1}{n}\sum_{i=1}^{n}\|\nabla h_{t}(\mathbf{X}_{T-t}^{i})\|_{2}^{2}\Bigg]$
$\leq\sum_{k=1}^{d}{\mathbb{E}}\Bigg[\sup_{h_{t}\in{\mathscr{H}}_{t}}\|D_{k}h_{t}\|_{L^{2}(p_{T-t})}^{2}-\frac{1}{n}\sum_{i=1}^{n}(D_{k}h_{t}(\mathbf{X}_{T-t}^{i}))^{2}\Bigg]$
$\leq\sum_{k=1}^{d}\frac{8\bar{B}^{2}}{\sigma_{T-t}^{4}}\Big(\frac{\mathrm{VCdim}(D_{k}{\mathscr{H}}_{t})}{n\log^{-1}n}\Big)^{\frac{1}{2}}\leq\frac{8d\bar{B}^{2}}{\sigma_{T-t}^{4}}\Big(\frac{\mathrm{VCdim}(\nabla{\mathscr{H}}_{t})}{n\log^{-1}n}\Big)^{\frac{1}{2}},$
where the first inequality holds from the convexity of supremum and Jensen’s inequality, the second inequality invokes a similar argument as Lemma D.2, and the last inequality holds from the definition of $\mathrm{VCdim}(\nabla{\mathscr{H}}_{t})$. This completes the proof. ∎
@@
${\mathbb{E}}\Bigg[\sup_{h_{t}\in{\mathscr{H}}_{t}}\frac{1}{n}\sum_{i=1}^{n}h_{t}(\mathbf{X}_{T-t}^{i})(w(\mathbf{X}_{0}^{i})-{\mathbb{E}}[w(\mathbf{X}_{0}^{i})\mid\mathbf{X}_{T-t}^{i}])\mid\mathbf{X}_{T-t}^{1:n}\Bigg]\leq 16\bar{B}^{2}\Big(\frac{\mathrm{VCdim}({\mathscr{H}}_{t})}{n\log^{-1}n}\Big)^{\frac{1}{2}},$
@@
Proof of Lemma D.5.
Define a sequence of auxilary random variables
$\xi_{i}\coloneq w(\mathbf{X}_{0}^{i})-{\mathbb{E}}[w(\mathbf{X}_{0}^{i})\mid\mathbf{X}_{T-t}^{i}].$
It is apparent that ${\mathbb{E}}[\xi_{i}\mid\mathbf{X}_{T-t}^{i}]=0$, and $|\xi_{i}|\leq\bar{B}$. Using Lemma D.4 yields
${\mathbb{E}}\Bigg[\sup_{h_{t}\in{\mathscr{H}}_{t}}\frac{1}{n}\sum_{i=1}^{n}\xi_{i}h_{t}(\mathbf{X}_{T-t}^{i})\mid\mathbf{X}_{T-t}^{1:n}\Bigg]$ $\leq 2\bar{B}{\mathbb{E}}\Bigg[\sup_{h_{t}\in{\mathscr{H}}_{t}}\frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}h_{t}(\mathbf{X}_{T-t}^{i})\mid\mathbf{X}_{T-t}^{1:n}\Bigg]$
$=2\bar{B}{\mathfrak{R}}({\mathscr{H}}_{t}\mid\mathbf{X}_{T-t}^{1:n}),$